I need pressure and Kp please...a The gas arsine, AsHz, decomposes as follows: 2AsH3(g) = 2As(s) + 3H2(g) In an experiment at a certain temperature, pure AsHCalculate K, for this reaction. K,=
Answer #1

Answer

Equilibrium pressure of H2 = 291 torr

Kp= 675.5

Explanation

2AsH3(g) <-----> 2As(s) + 3H2(g)

Kp = (PH2)3/(PAsH3)2

at equilibrium ,

PAsH3 = 385.0 - 2x

PH2 = 3x

After 48 hours reaction reaches equilibrium and the total pressure is 482.0 torr

(385.0 - 2x) + 3x = 482.0

x = 97 torr

PAsH3 = 385.0 - 2(97)= 191 torr

PH2 = 3 × 97 = 291 torr

Kp = ( 291 torr)3/(191torr)2

  Kp = 675.5